OpenStudy (anonymous):

Find the zeros of the function. f(x) = 4x^3 - 12x^2 - 40x

jimthompson5910 (jim_thompson5910):

f(x) = 4x^3 - 12x^2 - 40x 0 = 4x^3 - 12x^2 - 40x 4x^3 - 12x^2 - 40x = 0 4x(x^2 - 3x - 10) = 0 what's the next step?

OpenStudy (anonymous):

Distribute?

jimthompson5910 (jim_thompson5910):

distributing will only get you back to 4x^3 - 12x^2 - 40x = 0 so you'll be going backwards

OpenStudy (anonymous):

Well I'm not sure then.

jimthompson5910 (jim_thompson5910):

you need to factor x^2 - 3x - 10

OpenStudy (rajee_sam):

now factor the quadratic part

OpenStudy (anonymous):

(x-5)(x+3)

jimthompson5910 (jim_thompson5910):

close but no

jimthompson5910 (jim_thompson5910):

-5 times 3 = -15 and not -10

OpenStudy (anonymous):

No, sorry I ment to put 2 it was a typo

jimthompson5910 (jim_thompson5910):

good, x^2 - 3x - 10 factors to (x-5)(x+2)

OpenStudy (anonymous):

The keys are next to each other lol

jimthompson5910 (jim_thompson5910):

lol that's ok, happens to me all the time too

OpenStudy (anonymous):

Thanks for the help:)

jimthompson5910 (jim_thompson5910):

f(x) = 4x^3 - 12x^2 - 40x 0 = 4x^3 - 12x^2 - 40x 4x^3 - 12x^2 - 40x = 0 4x(x^2 - 3x - 10) = 0 4x(x-5)(x+2) = 0 4x = 0 or x-5 = 0 or x+2 = 0 x = ?? or x = ?? or x = ??

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

@jim_thompson5910 , I have another queston, do you need a 0 in that too?

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

x=5 x=-2 do you need x=0?

jimthompson5910 (jim_thompson5910):

yes x = 0 is a solution there should be 3 solutions total

OpenStudy (anonymous):

Oh, okay. Thanks:)

jimthompson5910 (jim_thompson5910):

if you were to start with f(x) = 4x^3 - 12x^2 - 40x and plug in x = 0, you would get f(x) = 0 as a result

jimthompson5910 (jim_thompson5910):

so that proves x = 0 is a solution yw